a square has an area of 36 square centemeters a rectangle has the same area as the square if the rectangle long sides mesure 9 centemters what is the rectangles perimeter

Answer by ilana(307)

If the rectangle's area is 36 square centimeters, and its length is 9 cm, then its width must be 36/9 = 4 cm.
perimeter = 2l + 2w
= 2*9 + 2*4
= 18 + 8
= 26 cm